# ArtFormer: Controllable Generation of Diverse 3D Articulated Objects

### v1

@@ Abstract
Figure 1: We present the Articulation TransFormer, for high-quality generation articulated objects. This figure illustrates controlled generation across random trials based on text descriptions. The openness ratio is defined within the generated joint limits and sampled for visualization. Notably, it can generate a diverse range of objects with varying numbers of sub-parts and different geometry features.
@@ 1 Introduction
To achieve both diversity and usability, this paper proposes a novel framework, Articulation Transformer, to generate high-quality and diverse articulated objects from text descriptions. We parameterize each articulate object with a tree structure. Each node corresponds to a sub-part, encompassing both its geometry and the kinematic relation (joint transform) relative to its parent node. Treating each node as a token, we utilize a transformer architecture to generate the sub-parts of the articulated object. Additionally, we introduce a tree position embedding in place of the ordinary position embeddings to better encode the tree structure from a sequence of tokens. Conditions (such as text descriptions and images) can be flexibly incorporated using cross-attention modules in the transformer layers.
@@
In this paper, we primarily conduct experiments the on text-guided generation of articulated objects, with a pre-trained text-encoder providing conditions to the transformer. Compared to prior works, our results demonstrate that we can generate a more diverse array of articulate objects that exhibit more precise kinematic features and high-quality geometry as well. Moreover, we also validate the flexibility of our framework with image-guided generation.
@@ 2 Related Work
A-SDF [38] is among the earliest explorers of using SDF to model articulated objects, but did not utilize the aforementioned property. More recently, NAP [25] introduces the first 3D deep generative model for synthesizing articulated objects through a novel articulation tree/graph parameterization and the use of a DDPM [14], enabling masked generation applications. Similarly, CAGE [30] also employs a graph diffusion denoising model but with a primary aim of controllability. SINGAPO [28] further extends controllable generation to single-image conditioning.
@@ 2.2 Generating Tree-structured Objects
Generating tree-structured objects differs from conventional sequential generation, as each node can have multiple successors. Traditional approaches model graph distributions using Variational Autoencoders (VAEs) [24, 53, 63], Generative Adversarial Networks (GANs) [57, 8, 33] and Denoising Diffusion Networks [16, 17, 21, 65, 64]. DiGress [56] and FreeGress [39], for example, achieve state-of-the-art generation performance and can handle large, diverse molecular datasets. However, these methods are not tailored for tree-structured graphs and lack autoregressive generation, resulting in unreliable outputs for realistic acyclic, single-edge tree structures. To address this limitation, SceneHGN [11] introduces a recursive autoencoder-based method, enabling hierarchical, tree-structured generation of 3D indoor scenes. Similarly, ATISS [42] achieves hierarchical generation using a novel autoregressive transformer architecture that generates 3D indoor scenes as unordered object sets. Shiv et al. [51] extend transformers to tree-structured data by propose a novel tree-to-sequence mapping method. Peng et al. [44] advance this approach and enable Transformers to learn from both pairwise node paths and leaf-to-root paths by integrating tree path encoding into the attention module.
@@ 3.2 Diverse and Controllable Shape Prior
Shape Prior. An articulated object consists of multiple sub-parts. Given a sub-part sampled from the dataset, we encode its point cloud with a VAE encoder: $q(z|f)$ where $f=\Gamma(\text{pcd})\in\mathbb{R}^{3\times 256\times 64\times 64}$ is an intermediate tri-plane feature obtained from a PointNet encoder $\Gamma$. A generalizable SDF network $\Omega(f,x)$ then predicts the part’s SDF at query points $x\in R^{3}$ from decoded features $p(f|z)$. The training objective is:
@@
However, $p(z)$ To subsequently enable guided or conditional generation of object sub-parts geometry, we train a conditional diffusion model $\epsilon(z_{t},t,\hat{c}_{g},c_{s})$ on $p(z)$, where the geometry and semantic conditions are given by two corresponding encoders: $c_{g}=\mathcal{E}_{g}(z)$ and $c_{s}=\mathcal{E}_{s}(name)$. Before geometry condition $c_{g}$ is input into the diffusion model, it is processed into $\hat{c}_{g}$ using codebooks. The diffusion model is trained to denoise random latent $z_{T}\sim\mathcal{N}(0,\mathbf{I})$ into a meaningful $z_{0}\sim p(z)$, following the objective used in [48]:
@@
Sampling Diverse Shapes. A particularly desirable capability is to generate parts with diverse geometry features given its semantic information. For example, we would like USB caps of different shapes and styles. To enable our model for such capability, we discretize the space of geometry code $c_{g}=\mathcal{E}_{g}(z)$ to allow for sampling. A geometry condition $c_{g}$ is chunked into 4 segments $(c_{g}^{0},c_{g}^{1},c_{g}^{2},c_{g}^{3})$ and used to retrieve $(\hat{c}_{g}^{0},\hat{c}_{g}^{1},\hat{c}_{g}^{2},\hat{c}_{g}^{3})$ from 4 different codebooks $M_{i}\in\mathbb{R}^{N\times D}$ using Gumbel-Softmax sampling:
$\hat{c}_{g}^{i}=\sum_{j=1}^{N}\limits{m^{i}_{j}\cdot\operatorname{GS}\left(\{-||m^{i}_{l}-c_{g}^{i}||_{2}\}_{l=1}^{N}\right)_{j}},$ (3)
where $m^{i}_{l}\in\mathbb{R}^{D}$ denotes the $l$-th out of $N$ embedding vector in the codebook $M_{i}$. The Gumble-Softmax operation is defined as:
$\operatorname{GS}(\{x_{k}\})_{i}=\frac{\exp{\left((x_{i}+g_{i})/\tau\right)}}{\sum_{j=1}^{D}\limits{\exp{\left((x_{j}+g_{j})/\tau\right)}}},$ (4)
where $g_{1},\cdots,g_{k}$ are samples from $\operatorname{Gumbel}(0,1)$ [20]. The softmax temperature $\tau$ controls the diversity of shape prior, which we do not specifically tune in this work. Since $GS$ sampling is differentiable, the model can still be trained end-to-end.
Later, for the $i$-th sub-part, we directly predict $c_{s}$ and logits $P_{i}\in\mathbb{R}^{4\times N}$ to sample from the codebooks $M$ and form $\hat{c}_{g}$. We avoid directly predicting the high-dimensional $z$ when generating the geometry, which is a great challenge in terms of model capacity. Meanwhile, the stochasticity introduced by discrete sampling improves the generation diversity.
@@ 3.3 Articulation Transformer
After the articulated object is parameterized into a tree structure, each node is treated as a token in the classical transformer architecture. $i$-th token is composed of $[\text{fa}_{i},b_{i},z_{i},j_{i},l_{i}]$, where $\text{fa}_{i}$ is the parent index of $i$-th node. In addition to the index of the parent node $\text{fa}_{i}$, the remaining attributes are concatenated and subsequently processed through an MLP tokenizer. An overview of the articulation transformer is illustrated in Fig. 3.
Tree Position Embedding. In order for the transformer to recognize the specific position of each token, we proposed a novel position encoding scheme specifically designed for tree structures building upon the works of [52] and [45]. We first calculate the absolute position encoding $a_{i}$ for each $i$-th node:
$a_{i}=\operatorname{GRU}\left(\{\text{Path}_{k}\}_{k=\mathcal{R}}^{i}\right).$ (5)
It push the tokens on the path from the root $\mathcal{R}$ to the $i$-th node to a bi-directional GRU [45, 23] to compress the information on its path. We define the position embedding of the $i$-th node $p_{i}$ to represent the relative position as well:
$p_{i}=\operatorname{CAT}\left(\{a_{k}\}_{k=i}^{\mathcal{R}}\right),$ (6)
@@
The decoding procedure terminates when all output tokens are $\mathcal{T}$. We then un-tokenize the generated tokens to the format described in Sec. 3.1, i.e., the kinematic characteristics and relations $(b_{i},j_{i},l_{i})$, and conditions $(P,c_{s})_{i}$ for SDF decoding. To obtain the final object, cast joint transforms to rigid transforms to the input of SDF $\Omega$ and extract mesh using the Marching Cube [31] algorithm.
@@
The training of the shape prior is conducted first. Upon completion, for each sub-part’s geometry latent code $z$ in dataset, we compute a matrix $\mathcal{D}$:
$\mathcal{D}_{i,j}=-||m_{j}^{i}-c_{g}^{i}||_{2},$ (7)
where $(c_{g}^{0},c_{g}^{1},c_{g}^{2},c_{g}^{3})=c_{g}=\mathcal{E}_{g}(z)$ and $m_{j}^{i}$ denotes $j$-th embedding vector in the codebook $M_{i}$. $L_{P}$ denotes the loss function to supervise $P$:
@@ 4 Experiments
Figure 5: Qualitative comparison between ArtFormer and baselines. Our method is capable of generating high-quality geometry and accurate joint relations.
@@ 4.1 Experimental Setup
We train the shape prior on PartNet [36] and PartNet-Mobility [61]. Although PartNet does not provide kinematic information, it still contributes to learning the geometry. The ArtFormer and other baseline models are trained exclusively on PartNet-Mobility. For each articulated object in the dataset, we use Blender to create high-resolution thumbnails and employ ChatGPT-4o to generate corresponding descriptions, which are used for training the baselines. Detailed implementation steps can be found in the supplementary material.
@@ 4.3 Metrics
A human study methodology (HS) is used to assess the alignment between generated objects and their text descriptions, as well as the diversity of the generated objects. $P=20$ participants are presented with objects generated from $M=20$ distinct descriptions, each by 4 baseline models. Participants select the object that best matches the description. In a separate task, they choose the most diverse group from 4 options, each containing objects generated from the same instruction by different baselines. This task is repeated $T=5$ times. The alignment score (AL) and diversity score (DS) are defined as the mean win rate. Further details are provided in the Supplementary Material.
@@ 4.4 Results
Table 1: Comparison of Generation Quality
Part Retrieval $\text{POR}\downarrow$${}_{\times 10^{-2}}$ ID HS
MMD $\downarrow$ COV $\uparrow$ 1-NNA $\downarrow$ AL $\uparrow$ DS $\uparrow$
NAP-128 0.805 0.0710 0.3085 0.7021 0.105 0.13
NAP-768 1.620 0.0632 0.3723 0.6543 0.093 0.12
Ours 0.709 0.0292 0.5213 0.5266 0.459 0.67
CAGE 0.251 0.0193 0.6064 0.5319 0.343 0.07
Ours-PR 0.556 0.0214 0.6400 0.3950 - -
@@
Novel Shape Generation. We conducted an experiment inspired by Diffusion-SDF to demonstrate that our shape prior, guided by an articulation transformer, can generate new geometry shapes that never appear in the dataset. We used our model to produce various objects and dissected them into sub-parts. Then, we calculated the Chamfer Distance between each sub-part and those in the training set and ranked them from nearest to farthest. The results, shown in Fig. 6, indicate that the sub-parts generated by our model are distinct from those in the training set, confirming the model’s ability to create novel geometry shapes.
@@ 4.5 Ablations
In our ablation study (Tab. 2), we evaluated the impact of removing the tree position embedding (TPE) and shape prior (SP). Without the shape prior, the transformer must directly generate sub-part geometry, complicating feature generation, eliminating randomness, and significantly degrading all performance metrics. Removing the tree position embedding causes the model to lose positional information, increasing POR due to sub-part overlap during motion and reducing COV by impairing its ability to capture structural nuances and dataset distribution.
@@ 5 Conclusion
We propose a novel method for modeling and generating 3D articulated objects, addressing limitations in diversity and usability. Representing articulated objects as a tree structure with labeled nodes for rigid parts simplifies articulation parameterization, enabling part-level definition and generation. To ensure well-aligned yet diverse outputs, we develop a controllable shape prior and the an Articulation Transformer, which captures articulation features effectively. A tree position embedding layer enhances part relationship modeling, supporting autoregressive generation. Our framework achieves state-of-the-art performance, generating high-quality, diverse objects from text and image conditions.
Limitations and Future Work. Our method has several limitations that warrant further investigation: (1) The limited dataset restricts the application to a small range of object types with few components, preventing the full potential of the approach from being realized. Owing to the transformer backbone, we posit that our architecture is more adept at generating long sequences, such as objects with numerous sub-parts. Future work may explore this capability in more detail. (2) Multi-modal instructions beyond text and images have not yet been explored, such as point cloud or joint structure of expected articulated object. Investigating diverse instruction formats could greatly enhance flexibility and usability of our method in application. (3) Capturing articulation details in the text condition, such as rotation angles, is more challenging than category and geometry condition. Further research is needed to improve representation and learning of this data.

### v2

@@ Abstract
github.com/ShuYuMo2003/ArtFormer
For the convenience of readers, the changes made during the rebuttal stage are highlighted in different colors. [Uncaptioned image]
Figure 1: We present the Articulation TransFormer, for high-quality generation articulated objects. This figure illustrates controlled generation across random trials based on text descriptions. Notably, it can generate a diverse range of objects with varying numbers of sub-parts and different geometry features.
@@ 1 Introduction
To achieve both diversity and usability, this paper proposes a novel framework, Articulation Transformer, to generate high-quality and diverse articulated objects from text descriptions. We parameterize each articulated object with a tree structure. Each node corresponds to a sub-part, encompassing both its geometry and the kinematic relation (joint transform) relative to its parent node. Treating each node as a token, we utilize a transformer architecture to generate the sub-parts of the articulated object. Additionally, we introduce a tree position embedding in place of the ordinary position embeddings to better encode the tree structure from a sequence of tokens. Conditions (such as text descriptions and images) can be flexibly incorporated using cross-attention modules in the transformer layers.
@@
In this paper, we primarily conduct experiments on text-guided generation of articulated objects, with a pre-trained text-encoder providing conditions to the transformer. Compared to prior works, our results demonstrate that we can generate a more diverse array of articulated objects that exhibit more precise kinematic features and high-quality geometry as well. Moreover, we also validate the flexibility of our framework with image-guided generation.
@@ 2 Related Work
A-SDF [42] is among the earliest explorers of using SDF to model articulated objects, but did not utilize the aforementioned property. More recently, NAP [28] introduces the first 3D deep generative model for synthesizing articulated objects through a novel articulation tree/graph parameterization and the use of a DDPM [16], enabling masked generation applications. Similarly, CAGE [34] also employs a graph diffusion denoising model but with a primary aim of controllability. SINGAPO [32] further extends controllable generation to single-image conditioning. MeshArt [12] utilizes domain-specific tokenizers to convert articulated objects into sequences and employs a transformer architecture for generation.
@@ 2.2 Generating Tree-structured Objects
Generating tree-structured objects differs from conventional sequential generation, as each node can have multiple successors. Traditional approaches model graph distributions using Variational Autoencoders (VAEs) [27, 57, 67], Generative Adversarial Networks (GANs) [61, 9, 37] and Denoising Diffusion Networks [18, 19, 23, 69, 68]. DiGress [60] and FreeGress [43], for example, achieve state-of-the-art generation performance and can handle large, diverse molecular datasets. However, these methods are not tailored for tree-structured graphs and lack autoregressive generation, resulting in unreliable outputs for realistic acyclic, single-edge tree structures. To address this limitation, SceneHGN [13] introduces a recursive auto-encoder-based method that enables the hierarchical tree-structured generation of 3D indoor scenes. Similarly, GRAINS [31] achieves hierarchical indoor scenes generation using a specific recursive VAE. Shiv et al. [55] extend transformers to tree-structured data by proposing a novel tree-to-sequence mapping method. Peng et al. [48] advance this approach and enable Transformers to learn from both pairwise node paths and leaf-to-root paths by integrating tree path encoding into the attention module.
@@ 3.2 Diverse and Controllable Shape Prior
Shape Prior. An articulated object consists of multiple sub-parts. Given a sub-part sampled from the dataset, we encode its point cloud with a VAE encoder: $q(z|f)$ where $f=\Gamma({\color[rgb]{1,0,1}\text{point cloud}})\in\mathbb{R}^{3\times 256\times 64\times 64}$ is an intermediate tri-plane feature obtained from a PointNet encoder $\Gamma$. A generalizable SDF network $\Omega(f,x)$ then predicts the part’s SDF at query points $x\in R^{3}$ from decoded features $p(f|z)$. The training objective is:
@@
However, to subsequently enable guided or conditional generation of object sub-parts geometry, we train a conditional diffusion model $\epsilon(z_{t},t,\hat{c}_{g},c_{s})$ on $p(z)$, where the geometry and semantic conditions are given by two corresponding encoders: $c_{g}=\mathcal{E}_{g}(z)$ and $c_{s}=\mathcal{E}_{s}(name)$. Before geometry condition $c_{g}$ is input into the diffusion model, it is processed into $\hat{c}_{g}$ using codebooks. The diffusion model is trained to denoise random latent $z_{T}\sim\mathcal{N}(0,\mathbf{I})$ into a meaningful $z_{0}\sim p(z)$, following the objective used in [52]:
@@
Sampling Diverse Shapes. A particularly desirable capability is to generate parts with diverse geometry features given its semantic information. For example, we would like USB caps of different shapes and styles. To enable our model for such capability, we discretize the space of geometry code $c_{g}=\mathcal{E}_{g}(z)$ to allow for sampling. A geometry condition $c_{g}$ is chunked into 4 segments $(c_{g}^{0},c_{g}^{1},c_{g}^{2},c_{g}^{3})$, which is designed to enlarge the capacity of the latent space from $4N$ to $N^{4}$. Then, these 4 segments are used to retrieve $(\hat{c}_{g}^{0},\hat{c}_{g}^{1},\hat{c}_{g}^{2},\hat{c}_{g}^{3})$ from 4 different codebooks $M_{\color[rgb]{1,0,1}t}\in\mathbb{R}^{N\times D}$ ($t$ is the index of codebook) using Gumbel-Softmax sampling:
$\mathcal{D}_{t,l}=-||m_{l}^{{\color[rgb]{1,0,1}t}}-c_{g}^{\color[rgb]{1,0,1}t}||_{2}$ (3)
$\hat{c}_{g}^{\color[rgb]{1,0,1}t}=\sum_{j=1}^{N}\limits{m_{j}^{{\color[rgb]{1,0,1}t}}\cdot\operatorname{GS}\left(\{\mathcal{D}_{t,l}\}_{l=1}^{N}\right)_{j}}$ (4)
where $\mathcal{D}_{t,l}$ is the distance matrix and $m_{l}^{{\color[rgb]{1,0,1}t}}\in\mathbb{R}^{D}$ denotes the $l$-th out of $N$ embedding vector in the codebook $M_{\color[rgb]{1,0,1}t}$. The Gumble-Softmax operation is defined as:
$\operatorname{GS}\left(\{x_{\color[rgb]{1,0,1}l}\}_{{\color[rgb]{1,0,1}l=1}}^{\color[rgb]{1,0,1}N}\right)_{j}=\frac{\exp{\left((x_{j}+g_{j})/\tau\right)}}{\sum_{{\color[rgb]{1,0,1}k}=1}^{{\color[rgb]{1,0,1}N}}\limits{\exp{\left((x_{\color[rgb]{1,0,1}k}+g_{\color[rgb]{1,0,1}k})/\tau\right)}}}$ (5)
where $g_{1},\cdots,g_{k}$ are samples from $\operatorname{Gumbel}(0,1)$ [22]. The softmax temperature $\tau$ controls the diversity of shape prior, which we do not specifically tune in this work. Since $\operatorname{GS}$ sampling is differentiable, the model can still be trained end-to-end.
The diffusion model helps sample from the full continuous space of $z$, given $\hat{c}_{g}$ which is either quantised from $c_{g}$ (produced by $\mathcal{E}_{g}$, during shape prior training) or sample by logits $P_{i}\in\mathbb{R}^{4\times N}$ from codebook $M$ directly (after shape prior training). The shape prior ensure the sampled $z$ is well in distribution. Meanwhile, the stochasticity introduced by discrete sampling improves the generation diversity.
@@ 3.3 Articulation Transformer
After the articulated object is parameterized into a tree structure, each node is treated as a token in the classical transformer architecture. $i$-th token is composed of $[\text{fa}_{i},b_{i},z_{i},j_{i},l_{i}]$, where $\text{fa}_{i}$ is the parent index of $i$-th node. In addition to the index of the parent node $\text{fa}_{i}$, the remaining attributes are concatenated and subsequently processed through an MLP Embedding Mapper. An overview of the articulation transformer is illustrated in Fig. 3.
Tree Position Embedding. In order for the transformer to recognize the specific position of each token, we proposed a novel position encoding scheme specifically designed for tree structures building upon the works of [56, 49, 25]. We first calculate the absolute position encoding $a_{i}{\color[rgb]{1,0,1}\in\mathbb{R}^{64}}$ for each $i$-th node following [49]:
$a_{i}=\operatorname{GRU}\left(\{{\color[rgb]{1,0,1}\mathcal{A}_{p_{k}}}\}_{k=1}^{K}\right)$ (6)
where $\mathcal{A}_{i}=[b_{i},z_{i},j_{i},l_{i}]$, $K$ is the depth of $i$-th node and $\{p_{k}\}_{k=1}^{K}$ is the sequence of index alone the path on the tree from the root to $i$-th node. It push the tokens on the path from the root to a bi-directional GRU [49, 25] to compress the information. We define the position embedding of the $i$-th node $p_{i}{\color[rgb]{1,0,1}\in\mathbb{R}^{1024}}$ to represent the relative position following [56]:
$p_{i}=\operatorname{CAT}\left(\{a_{{\color[rgb]{1,0,1}p_{k}}}\}_{k=K}^{{\color[rgb]{1,0,1}1}}\right),$ (7)
@@
The decoding procedure terminates when all output tokens are $\mathcal{T}$. We then project the generated tokens back to the format described in Sec. 3.1, i.e., the kinematic characteristics and relations $(b_{i},j_{i},l_{i})$, and conditions $(P,c_{s})_{i}$ for SDF decoding. To obtain the final object, cast joint transforms to rigid transforms to the input of SDF $\Omega$ and extract mesh using the Marching Cube [35] algorithm.
@@
The training of the shape prior is conducted first. Upon completion, for each sub-part’s geometry latent code $z$ in dataset, we compute the matrix $\mathcal{D}$ defined in  Eq. 3.
$L_{P}$ denotes the loss function to supervise $P$:
@@ 4 Experiments
Table 1: Comparison of Generation Quality
Part Retrieval $\text{POR}\downarrow$${}_{\times 10^{-2}}$ ID HS
MMD $\downarrow$ COV $\uparrow$ 1-NNA $\downarrow$ AL $\uparrow$ DS $\uparrow$
NAP-128 0.805 0.0710 0.3085 0.7021 0.105 0.13
NAP-768 1.620 0.0632 0.3723 0.6543 0.093 0.12
Ours-NAPSP 2.761 0.0375 0.4831 0.8315 - -
Ours 0.709 0.0292 0.5213 0.5266 0.459 0.67
CAGE 0.251 0.0193 0.6064 0.5319 0.343 0.07
Ours-PR 0.556 0.0214 0.6400 0.3950 - -
Figure 5: Qualitative comparison between ArtFormer and baselines (Ours-1CB will be discussed in Sec. 4.5). Our method is capable of generating high-quality geometry and accurate joint relations.
@@ 4.1 Experimental Setup
We train the shape prior on PartNet [40] and PartNet-Mobility [65]. Although PartNet does not provide kinematic information, it still contributes to learning the geometry. The ArtFormer and other baseline models are trained exclusively on 6 categories (Storage Furniture, Safe, Oven, USB, Bottle and Washer) in PartNet-Mobility. For each articulated object in the dataset, we use Blender to create high-resolution thumbnails and employ ChatGPT-4o to generate corresponding descriptions, which are used for training the baselines. Detailed implementation steps can be found in Supplementary Material.
@@ 4.2 Baselines
  6. 6.
    Ours-NAPSP: Building on our original model, but the geometry is generated through the shape prior of NAP.
@@ 4.3 Metrics
A human study methodology (HS) is used to assess the alignment between generated objects and their text descriptions, as well as the diversity of the generated objects. $P=20$ participants are presented with objects generated from $M=20$ distinct descriptions, each by 4 baseline models. Participants select the object that best matches the description. In a separate task, they choose the most diverse group from 4 options, each containing objects generated from the same instruction by different baselines. This task is repeated $T=5$ times. The alignment score (AL) and diversity score (DS) are defined as the mean win rate. Further details are provided in Supplementary Material.
@@ 4.5 Ablations
In our ablation study (Tab. 2), we evaluated the impact of removing the tree position embedding (TPE) and shape prior (SP). Without the shape prior, the transformer must directly generate sub-part geometry, complicating feature generation, eliminating randomness, and significantly degrading all performance metrics. Removing the tree position embedding causes the model to lose positional information, increasing POR due to sub-part overlap during motion and reducing COV by impairing its ability to capture structural nuances and dataset distribution. We conduct the experiment with a single codebook (1 Codebook), and the performance drop is shown in Fig. 5 and Tab. 2.
@@ 5 Conclusion
We propose a novel method for modeling and generating 3D articulated objects, addressing limitations in diversity and usability. Representing articulated objects as a tree structure with nodes for rigid parts simplifies articulation parameterization, enabling part-level definition and generation. To ensure well-aligned yet diverse outputs, we develop a controllable shape prior and the an Articulation Transformer, which captures articulation features effectively. A tree position embedding layer enhances part relationship modeling, supporting autoregressive generation. Our method achieves state-of-the-art performance, generating high-quality, diverse objects from text or image conditions.
Limitations and Future Work. (1) The limited dataset restricts the application to a small range of object types with few sub-part, preventing the full potential of the approach from being realized. Our model is expected to show more advantage when the number of parts is large ($>10$) and varies (which are rare in existing dataset), due to transformer’s capability to handle long sequences of varying lengths. Future work may explore this capability in large scale. (2) Multi-modal instructions beyond text and images have not yet been explored, such as point cloud or joint structure of expected articulated object. Investigating diverse instruction formats could greatly enhance flexibility and usability of our method in application. (3) Capturing joint quantitative details in the text condition, such as rotation angles, is more challenging than joint type and geometry condition. Further research is needed to improve representation and learning of this data. (4) It is observed that increasing the number of object categories during training leads to a decline in geometry reconstruction quality, likely due to the limited generalization of SDF models. Methods from GenSDF [5] may enhance generalization, and we also recommend using modern 3D representations like 3D Gaussian Splatting [26] in future work.
Acknowledgements. This paper is a pure student work, as all authors are either undergraduate or PhD students. The computational resources were provided by Ningxia Westcloud Computing Technology Co., Ltd. (Damodel), which offered us certain discounts and waivers. Partial financial support was provided by Xiamen University Malaysia.
@@
For the convenience of readers, the changes made during the rebuttal stage are highlighted in different colors.